OpenStudy (waheguru):

Find the x value y=x+b 2=x+4 I am getting -2 but the answer is 2

OpenStudy (waheguru):

2-4=x -2=x?

OpenStudy (anonymous):

-2 is the correct answer.... UNLESS the question is written down incorrectly....